Parth (parthkohli):

Condition on \(a,b\) that if \(\alpha\) is a root of \(x^2 + ax + b\), then so is \(\alpha^2 -2\). Hold on for my attempt...

Parth (parthkohli):

\[\alpha^2 + a\alpha + b = 0\]So I divide \((\alpha^2 -2)^2 + a (\alpha^2 - 2) + b \) by \(\alpha^2 + a\alpha + b\) and equate the dividend with the remainder. Correct?

Parth (parthkohli):

Well, long division is turning out to be really long.

hartnn (hartnn):

how about using the fact that product of roots is b and sum is -a

Parth (parthkohli):

\[c + c^2 -2 = -a\Rightarrow c^2 + c + a-2 = 0\]\[c(c^2 - 2) = b \Rightarrow c^3 - 2c - b = 0\]In the first equation, \(1 - 4(a - 2) > 0\). In the second equation... well...

Parth (parthkohli):

I find the long division method better - but am too lazy to perform it. :P

Parth (parthkohli):

Or maybe I should solve the first equation and plug that into the second.

Parth (parthkohli):

\[c = \dfrac{-1 \pm \sqrt{1 - 4(a-2)}}{2} = \dfrac{-1 \pm \sqrt{4a - 9}}{2}\]

Parth (parthkohli):

Whoops.

Parth (parthkohli):

The radical should contain \(9 - 4a.\)

Parth (parthkohli):

Not that it makes the problem any simpler...

OpenStudy (asnaseer):

@ParthKohli - surely what @hartnn said is the simplest way to get the conditions on a and b?

OpenStudy (asnaseer):

is it sufficient to express both a and b in terms of \(\alpha\)?

Parth (parthkohli):

I don't know. What I do know that the number of values of \(a,b\) taken together would be finite.

OpenStudy (asnaseer):

how?

Parth (parthkohli):

Because that's what the question is. It asks me for the number of values of a,b taken together.

Parth (parthkohli):

For that, I'd have to get a relation between a and b - I suppose.

OpenStudy (asnaseer):

Making use of what @hartnn said we get:\[a=2-\alpha-\alpha^2\]\[b=\alpha(\alpha^2-2)\]you can set \(\alpha\) to any integer between \(-\infty\) and \(\infty\) and it will generate an infinite number of a's and b's - wouldn't it?

OpenStudy (asnaseer):

unless there is some other condition like both a and b have to be positive integers, etc?

Parth (parthkohli):

Nope - real.

OpenStudy (asnaseer):

then I cannot see how the number of solutions to this is finite?

Parth (parthkohli):

There must be another restricting factor. :|

Parth (parthkohli):

You're the smart guy in the room - you should tell me. x)

Parth (parthkohli):

@ganeshie8

OpenStudy (anonymous):

is it that ax^2+bx+c=0 has its two roots as alpha and alpha^2-2

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

\[(x-\alpha)(x-(\alpha^2-2))=0 and ax^2+bx+c=0\] represents same

Parth (parthkohli):

\[a(x- \alpha)(x - \alpha^2 + 2) = ax^2 + bx + c\]But again, that is almost the same as hartnn's method.

OpenStudy (anonymous):

 by comparing co efficients